There are different configurations of hybrid electric vehicles (HEVs). On of its configuration is power split hybrid electric vehicle. One of its applications can be found out in Toyota Prius that utilized Hybrid Synergy Drive THS-II. A key component of power split hybrid electric vehicle is planetary gear which is responsible for combining Internal Combustion Engine (ICE), Electric Motor and Generator.


You are required to describe its kinematic. Also solve for its kinetics if the following information are provided.

The desired torque of vehicle = 100+last two digits of your registration [Nm] Mass of vehicle = 1504 kg

Final gear ratio from wheel to the ring = 3.93 Torques of Electric motor = 200 [Nm] Torques of ICE = 90 [Nm]

 

speed of Electric motor = 500 [rad/s] speed of ICE = 400 [rad/s]

the moment of inertia of engine, motor, generator, and vehicle are 1.746 [kgm2], 0.0226 [kgm2], 0.0226 [kgm2], and 3.3807 [kgm2] respectively.

Find toque and speed of generator.


1
Expert's answer
2022-02-07T23:44:02-0500

hybrid electric vehicles (HEVs) are powered by an internal combustion engine in combination with one or more electric motors that use energy stored in batteries. HEVs combine the benefits of high fuel economy and low tailpipe emissions with the power and range of conventional vehicles.

"Power = Torque \u00d7 rotational speed(rpm)"

Rpm = rotation per 1 minute.


unit of rotational speed as rad/s (radian per second)

1 rotation = 360 degree = 6.28 rad


So "1 rpm = \frac{6.28}{60} \frac{rad}{sec}.= 0.10466667 \frac{rad}{s}"


unit of torque whether it is Nm or not.


"1Nm × 1\frac{rad}{s} = 1 W(watt)"


"P = 500 rpm × 200Nm = 500\frac{rad}{s} ×200Nm =100kW"